the perimeter of a rectangle is 600. The lenght of the rectangle is 2 more than 3 times the width.
Answer by nerdybill(7384)

the perimeter of a rectangle is 600. The lenght of the rectangle is 2 more than 3 times the width.
.
Let w = width of rectangle
then
3w+2 = length of rectangle
.
perimeter = 2(width+length)
.
600 = 2(w + 3w+2)
300 = w + 3w+2
300 = 4w+2
298 = 4w
74.5 = w (width)
.
length:
3w+2 = 3(74.5)+2 = 223.5+2 = 225.5 (length)